# VisionReasoner: Unified Visual Perception and Reasoning via Reinforcement Learning

### v3

@@ 4.4 Ablation Study
Reasoning Length. As shown in Table 5, our analysis reveals that the model’s reasoning length adapts dynamically to text query complexity. Specifically, for simple class names in COCO and short phrases in RefCOCOg, the reasoning process is relatively concise. In contrast, complex reasoning-intensive queries in ReasonSeg require longer reasoning processes.
@@
Multi-object Matching. We quantitatively assess the efficiency of our two key design choices for multi-object matching: the Hungarian algorithm and batch computation. In a scenario with 30 objects, Table 5 demonstrates that random permutation and simple matching require over seconds (i.e., years) to complete, while our optimized approach achieves matching in just seconds - a times speedup.
@@
Non Repeat Reward. Figure 4 (a) presents the performance comparison with and without the non-repeat reward. Models are trained only on 2,000 samples from RefCOCOg. The model achieves better results when trained using the non-repeat reward. Additionally, model without non-repeat reward tends to generate longer reasoning processes, as shown in Figure 4 (b), and we observe repetitive reasoning patterns during inference.
@@
Different Training Data. We conduct an ablation study on different training datasets, with results presented in Section 4.3. The four datasets provide diverse text annotations: LVIS uses simple class names, RefCOCOg contains single-object referring expressions, gRefCOCO includes expressions that may refer to multiple objects, and LISA++ features texts requiring reasoning. Our experiments demonstrate that these datasets consistently improve model performance.
@@
Visual QA Ability. We also compare VisionReasoner’s VQA [26, 27, 23, 51, 46] ability with Qwen2VL [45] and the baseline model Qwen2.5VL[1]. As shown in Section 4.3, VisionReasoner achieves a slight performance gain even thought we do not train on VQA data.
@@
Sampling Number. Figure 5 presents performance comparison with different sampling number. Models are trained using all 7,000 training samples. We observe an initial performance gain followed by a notable decline with larger sampling number, suggesting that excessive sampling may induce overfitting to the training distribution and consequently degrade generalization capability.
@@
Reasoning. Figure 6 compares the performance of models with and without reasoning capabilities. We implement no reasoning by removing thinking reward. Models are trained using all 7,000 training samples. Our results show that both approaches outperform the baseline. And the reasoning-enhanced model demonstrates significant gain on intricate reasoning segmentation data.
@@ 4.5 Qualitative Results
We visualize some results on Figure 7. Notably, VisionReasoner addresses several visual perception tasks within a shared model. Our model generates comprehensive reasoning processes for all tasks while producing expected outputs. We find that VisionReasoner can effectively distinguish between similar objects, as shown in the visual grounding and referring expression segmentation. VisionReasoner also accurately localize multiple targets, as shown in object detection and counting. We also observe that the length of the reasoning process adapts dynamically: more intricate image-query pairs elicit detailed rationales, while simpler inputs result in concise explanations.
@@ 5 Conclusion
In this paper, we present VisionReasoner, a unified vision-language framework. With systematic task reformulation, VisionReasoner effectively addresses diverse visual perception tasks within a shared model. By introducing novel multi-object cognitive learning strategies and curated reward functions, VisionReasoner demonstrates strong capabilities in analyzing visual inputs, generating structured reasoning processes and delivering task-specific outputs. Experiments across ten diverse tasks—spanning detection, segmentation, and counting—validates the robustness and versatility of our approach. Notably, VisionReasoner achieves significant improvements over existing models such as Qwen2.5VL, with relative performance gains of 29.1% on COCO (detection), 22.1% on ReasonSeg (segmentation), and 15.3% on CountBench (counting). These results highlight the potential of unified frameworks in advancing multiple visual perception tasks.
Appendix A User Prompt Template
@@
To guide the policy model toward generating desired outputs during exploration, we employ the user prompt template presented in Table A1. This prompt template is inspired by DeepSeek-R1-Zero [8] and Seg-Zero [24].
Appendix B Concepts Clarification
@@
We formally define the key terms used in this work. As illustrated in Table A2, our hierarchical task formulation adopts the COCO dataset [17] as a representative example:
@@
| Task Type | Task category listed in PapersWithCode [35] (e.g., object detection) |
| Task | Concrete task (e.g., COCO object detection) |
Appendix C Details of Task Reformulation
Within our framework, we categorize task types as illustrated in Table A3 and Table A4. These tables highlight our grouping of task types based on their similarities. It is important to note that although this taxonomy covers a broad range of task types, the current implementation of VisionReasoner is evaluated on only 10 representative tasks, with comprehensive evaluation of all task types reserved for future research.
@@ Appendix C Details of Task Reformulation
Appendix D Task Router
We train TaskRouter using the GRPO algorithm [42], providing reward signals exclusively upon correct task classification. We evaluate the effective of the TaskRouter and results are shown on Table A5. The task classification dataset is constructed from diverse visual perception datasets and AI-generated samples. For datasets that include textual instructions (e.g., RefCOCOg), we retain their original instructions and corresponding fundamental task categories.
| Model | Accuracy |
| Qwen2.5-1.5B | 46.3 |
| TaskRouter-1.5B | 99.1 |
Additionally, for each fundamental task type, we employ ChatGPT [30] to generate instructions and target categories. The final dataset comprises 20,000 training samples and 4,000 test samples. Although the state-of-the-art Qwen2.5 [47] demonstrates strong performance in instruction following and zero-shot task classification, its accuracy drops below 50% in our complex scenario. In contrast, our task router module, trained using reinforcement learning, achieves significantly better performance.

### v4

@@ 4.3 Ablation Study
Multi-object Matching. We quantitatively assess the efficiency of our two key design choices for multi-object matching: the Hungarian algorithm and batch computation. In a scenario with 30 objects, Table˜5 demonstrates that a non-batch matching require seconds to complete, while our optimized approach achieves matching in just seconds - a speedup.
Reasoning Length. As shown in Table˜5, our analysis reveals that the model’s reasoning length adapts dynamically to text query complexity. Specifically, for simple class names in COCO and short phrases in RefCOCOg, the reasoning process is relatively concise. In contrast, complex reasoning-intensive queries in ReasonSeg require longer reasoning processes.
@@
Non Repeat Reward. Figure˜4 (a) presents the performance comparison with and without the non-repeat reward. Models are trained only on 2,000 samples from RefCOCOg. The model achieves better results when trained using the non-repeat reward. Additionally, model without non-repeat reward tends to generate longer reasoning processes, as shown in Figure˜4 (b), and we observe repetitive reasoning patterns during inference.
Different RL Algorithm. We use different on-policy RL training algorithm: the GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025). Models are trained only on 2,000 samples from RefCOCOg. As shown in Table˜5, performance consistently improves across both algorithms, demonstrating that our training framework is both stable and generalizable.
@@
Different Training Data. We conduct an ablation study on different training datasets, with results presented in Section˜4.2. The four datasets provide diverse text annotations: LVIS uses simple class names, RefCOCOg contains single-object referring expressions, gRefCOCO includes expressions that may refer to multiple objects, and LISA++ features texts requiring reasoning. Our experiments demonstrate that these datasets consistently improve model performance.
@@
Visual QA Ability. We also compare VisionReasoner’s VQA (Masry et al., 2022; Mathew et al., 2021; Liu et al., 2024b; Yue et al., 2024; xAI, 2024) ability with Qwen2VL (Wang et al., 2024) and the baseline model Qwen2.5VL(Bai et al., 2025). As shown in Section˜4.2, VisionReasoner achieves a slight performance gain even though we do not train on VQA data.
@@
Sampling Number. Figure˜5 presents performance comparison with different sampling number. Models are trained using all 7,000 training samples. We observe an initial performance gain followed by a notable decline with larger sampling number, suggesting that excessive sampling may induce overfitting to the training distribution and consequently degrade generalization capability.
@@
Reasoning. Figure˜6 compares the performance of models with and without reasoning capabilities. We implement no reasoning by removing thinking reward. Models are trained using all 7,000 training samples. Our results show that both approaches outperform the baseline. And the reasoning-enhanced model demonstrates significant gain on intricate reasoning segmentation data.
4.4 Human Evaluation on Reasoning Process
| IoU_Range | Num | IC (%) | AC (%) |
| 0–0.25 | |||
| 0.25–0.50 | |||
| 0.50–0.75 | |||
| 0.75–1.00 | |||
| ALL |
We conduct human evaluations on ReasonSeg-val to assess both answer consistency and image consistency of the reasoning trace. Image consistency measures whether the reasoning trace accurately describes the visual content in the image. Answer consistency evaluates whether the objects and information mentioned in the reasoning trace are consistent with the final predicted output. These evaluations help assess the faithfulness and reliability of the model’s internal reasoning.
We evaluate the reasoning traces at different levels based on IoU (between prediction and ground-truth) values and results are shown in Table˜8. We find that the overall image consistency accuracy and answer consistency accuracy reach 97.0% and 90.5%, respectively. The majority of problematic reasoning traces are concentrated in cases where the IoU is below 0.25. These results demonstrate that the reasoning trace of VisionReasoner is accurate and well-grounded, even though the model was trained without human-annotated reasoning data.
@@ 4.5 Qualitative Results
We visualize some results on Figure˜7. Notably, VisionReasoner addresses several visual perception tasks within a shared model. More results and the reasoning process are provided in Appendix˜E.
@@ 5 Conclusion
We present VisionReasoner, a unified vision-language framework for reasoning visual perception tasks. By introducing novel multi-object cognitive learning strategies and curated reward functions, VisionReasoner demonstrates strong capabilities in analyzing visual inputs, generating structured reasoning processes and delivering task-specific outputs. Experiments across ten diverse tasks, spanning detection, segmentation and counting, validates the robustness and versatility of our approach. Notably, VisionReasoner achieves significant improvements over baseline, with relative performance gains of 29.1% on COCO (detection), 22.1% on ReasonSeg (segmentation), and 15.3% on CountBench (counting). Human evaluation further reveals the reasoning traces of VisionReasoner are well-grounded and faithful.
Appendix A Details of Evaluation Benchmarks
| Type | Data | # of samples |
| Det | COCO | 36,781 |
| RefCOCO | 5,786 | |
| RefCOCO+ | 5,060 | |
| RefCOCOg | 7,596 | |
| Seg | RefCOCO | 1,975 |
| RefCOCO+ | 1,975 | |
| RefCOCOg | 5,023 | |
| ReasonSeg | 979 | |
| Count | Pixmo-Count | 1,064 |
| CountBench | 504 | |
| SUM | 66,023 |
We use ten benchmarks to evaluate model performance across general vision perception tasks. Our evaluation includes three fundamental task types: detection, segmentation and counting. Specifially, we employ COCO (Lin et al., 2014) and RefCOCO(+/g) (Yu et al., 2016) for detection evaluation; RefCOCO(+/g) and ReasonSeg (Lai et al., 2024) for segmentation evaluation; PixMo-Count (Deitke et al., 2024) and CountBench (Paiss et al., 2023) for counting evaluation.
Annotation Preparation. To ensure consistency across all evaluation tasks, we standardize the evaluation data by converting all samples into a unified multi-modal conversation format and removing potential information leakage. This preprocessing involves: converting numeric class labels to textual descriptions in COCO (Lin et al., 2014); removing explicit numerical references from text descriptions in CountBench (Paiss et al., 2023); applying consistent formatting across all datasets to maintain evaluation fairness.
Evaluation Metrics. For object detection on COCO, we adopt the standard AP metric computed using the COCO API (Team, 2014). For referring object grounding on RefCOCO(+/g), we use bbox AP, which measures detection accuracy at an IoU threshold of 0.5. For object segmentation on RefCOCO(+/g) and ReasonSeg, we use gIoU, computed as the mean IoU across all segmentation masks. For counting tasks, we use count accuracy as evaluation metric.
Statistics and Visualization. We show the statistic data in Table˜9. For detection and segmentation tasks, we report the number of valid instances. For counting tasks, we provide the total number of test samples. Our evaluation comprises a total of 66,023 test samples, covering three fundamental visual perception task types and 10 specific tasks. We visualize some examples in Figure˜8.
Appendix B Experission Level Evaluation on RefCOCO(+/g)
Our primary evaluation, detailed in Appendix˜A, reports instance-level performance. However, since the RefCOCO(+/g) benchmarks provide multiple expressions per image, we additionally present expression-level results in Table˜10.
| Method | Detection | SUM | |||||
| RefCOCO | RefCOCO+ | RefCOCOg | |||||
| val | testA | val | testA | val | test | ||
| Qwen2.5-VL-7B* | 90.0 | 92.5 | 84.2 | 89.1 | 87.2 | 87.2 | 530.2 |
| Qwen2.5-VL-7B | 89.0 | 92.0 | 83.2 | 88.3 | 86.4 | 86.5 | 525.4 |
[…]
Appendix C Task Router
We train TaskRouter using the GRPO algorithm (Shao et al., 2024), providing reward signals exclusively upon correct task classification. We evaluate the effective of the TaskRouter and results are shown on Table˜11. The task classification dataset is constructed from diverse visual perception datasets and AI-generated samples. For datasets that include textual instructions (e.g., RefCOCOg), we retain their original instructions and corresponding fundamental task categories.
| Model | Accuracy |
| Qwen2.5-1.5B | 46.3 |
[…]
Additionally, for each fundamental task type, we employ ChatGPT (OpenAI, 2023) to generate instructions and target categories. The final dataset comprises 20,000 training samples and 4,000 test samples. Although the state-of-the-art Qwen2.5 (Yang et al., 2024) demonstrates strong performance in instruction following and zero-shot task classification, its accuracy drops below 50% in our complex scenario. In contrast, our task router module, trained using reinforcement learning, achieves significantly better performance.
Appendix D User Prompt Template
@@
To guide the policy model toward generating desired outputs during exploration, we employ the user prompt template presented in Table˜12. This prompt template is inspired by DeepSeek-R1-Zero (Guo et al., 2025) and Seg-Zero (Liu et al., 2025a).
Appendix E Qualitative Results
We visualize results on Figure˜9. Our model generates comprehensive reasoning processes for all tasks while producing expected outputs. We find that VisionReasoner can effectively distinguish between similar objects, as shown in the visual grounding and referring expression segmentation. VisionReasoner also accurately localize multiple targets, as shown in object detection and counting. We also observe that the length of the reasoning process adapts dynamically: more intricate image-query pairs elicit detailed rationales, while simpler inputs result in concise explanations.
Appendix F Extension
More Applications. Our VisionReasoner can be easily extended to other visual perception tasks that need reasoning. We just need to add a light-weight module for different output format. The intermediate output and serve as bridge to connect other modules. Figure˜10 shows our extension to referring expression pose estimation and referring expression depth estimation.
Hybrid Mode. We can also employ a hybrid mode, that is directly using traditional visual models (e.g. Yolo-World (Cheng et al., 2024)) for simple categorical instruction (i.e. bird) and VisionReasoner for complex instructions (i.e. ‘Where can I have a rest?’).
Appendix G Concepts Clarification
@@
We formally define the key terms used in this work. As illustrated in Table 13, our hierarchical task formulation adopts the COCO dataset (Lin et al., 2014) as a representative example:
@@
| Task Type | Task category (e.g., object detection) |
| Task | Concrete benchmark (e.g., COCO object detection) |
Appendix H Details of Task Reformulation
Within our framework, we categorize task types as illustrated in Table˜14 and Table˜15. These tables highlight our grouping of task types based on their similarities. It is important to note that although this taxonomy covers a broad range of task types, the current implementation of VisionReasoner is evaluated on only 10 representative tasks, with comprehensive evaluation of all task types reserved for future research.